HV problems

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Tue Aug 28 2001 - 20:45:49 EDT

  • Next message: Apache: "Shift report 20010829 08:00-16:00"

    Dear Collaborator,
    
    We are again experience problems with the 4032 HV system that daisy chains 6 essential HV systems for BB, Mult , Gating grids,
    and TOFW calibration slats. These stopped working around noon today, and despite numerous efforts the crate could not cycled into operation. I therefore requested an emergency access period of 3-4 hours which will take place tomorrow. There is no guarentee we will identify the problem
    isolate and replace the faulty modules. Pending investigations it may be necessary to disbale parts of the experiment until we getr a long access period. This is unfortunate since the up-coming weekend + laboour day vacation has been planned as a long period for physics running without
    development work.
    On the positive front progress has been made on setting up the ZDC without saturations, and a short run to-day confirmed that we properly
    observe the 1n, 2n peaks in the spectrum.
    It is increasingly clear that problems develops slowly, so it is of uttermost importance that the shift people monitor scalers, rates, spectra to catch problems early. Please help out in this effort.
    
    Flemming
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Tue Aug 28 2001 - 20:37:12 EDT